OpenStudy (anonymous):

The length of a rectangle is twice its width. The perimeter of the rectangle is 90 cm. What are the length and width?

OpenStudy (anonymous):

\[l=2w\\ p=2(l+w)\qquad\text{=perimeter}\]

OpenStudy (anonymous):

so, we plug in \[90=2(2w+w)=2(3w)\\ 90=6w\\ w=?\]

OpenStudy (anonymous):

ok so how do i figure out what w is??

OpenStudy (anonymous):

90=6w divide both sides by 6

OpenStudy (anonymous):

15 and 1 ?